Consider the folloing data for two independent random samples taken from two normal populations with equal variances. find the 80% confidence interval for µ1 - µ2. sample 1: 12,8,11,6,13,7 sample 2: 13,16,10,9,13,14 what is the left endpoint and right endpoint?
Answer #1

Confidence Interval for Difference in two Population means \small \mu _{1}-\mu _{2}

Formula for Confidence Interval for Difference in two Population means when population Standard deviation are equal and not known

\small (\overline{x_{1}}-\overline{x_{2}})\pm t_{\alpha /2,n-1}\times s_{p}\sqrt{\frac{1}{n_{1}}+\frac{1}{n_{2}}}

\small Pooled\; \; Sample\; \; Standard\; deviation\; \; :s_{p}=\sqrt{\frac{(n_{1}-1)s_{1}^{2}+(n_{2}-1)s_{2}^{2}}{n_{1}+n_{2}-2}}

\small n_{1} : Sample Size of Sample 1
\small n_{2} : Sample Size of Sample 2
\small \overline{x_{1}} : Sample Mean of Sample 1
\small \overline{x_{2}} Sample Mean of Sample 2
s1 : Sample Standard Deviation of Sample 1
s2 : Sample Standard Deviation of Sample 2

For Sample 1

Sample 1 x1 x1-\small \overline{x_{1}} (x1-\small \overline{x_{1}})2
12 2.5 6.25
8 -1.5 2.25
11 1.5 2.25
6 -3.5 12.25
13 3.5 12.25
7 -2.5 6.25
Total 57 41.5
Mean: \small \overline{x_{1}} : 57/6=9.5

\small s_{1}=\sqrt{\sum \frac{(x_{1}-\overline{x}_{1})^{2}}{n_{1}-1}}=\sqrt{\frac{41.5}{6-1}}=2.881

For Sample 2

Sample 2 x2 x2-\small \overline{x}_{2} (x2-\small \overline{x}_{2})
13 0.5 0.25
16 3.5 12.25
10 -2.5 6.25
9 -3.5 12.25
13 0.5 0.25
14 1.5 2.25
Total 75 33.5
Mean 75/6=12.5

\small s_{2}=\sqrt{\sum \frac{(x_{1}-\overline{x}_{1})^{2}}{n_{1}-1}}=\sqrt{\frac{33.5}{6-1}}=2.588

n1 : Sample Size of Sample 1 6
n2 : Sample Size of Sample 2 6
\small \overline{x_{1}} : Sample Mean of Sample 1 9.5
\small \overline{x_{2}} : Sample Mean of Sample 2 12.5
s1 : Sample Standard Deviation of Sample 1 2.881
s2 : Sample Standard Deviation of Sample 2 2.588
Confidence Level 80%
\small \alpha (= 100-80/100=20/100 ) = 0.2 0.2
\small \alpha/2 (=0.2/2=0.1) 0.1
Degrees of freedom : ( n1+n2-2=6+6-2=10) 10
\small t_{\alpha /2,df}=t_{0.1,10} 1.3722

\small Pooled\; \; Sample\; \; Standard\; deviation\; \; :s_{p}=\sqrt{\frac{(n_{1}-1)s_{1}^{2}+(n_{2}-1)s_{2}^{2}}{n_{1}+n_{2}-2}}

\small s_{p}= \sqrt{\frac{(6-1)2.881^{2}+(6-1)2.588^{2}}{6+6-2}}=\sqrt{\frac{74.9895}{10}}=\sqrt{7.499}=2.7384

Confidence Interval for Difference in two Population means

\small (\overline{x_{1}}-\overline{x_{2}})\pm t_{\alpha /2,n-1}\times s_{p}\sqrt{\frac{1}{n_{1}}+\frac{1}{n_{2}}}

\small =(9.5-12.5)\pm1.3722\times2.7384 \sqrt{\frac{1}{6}+\frac{1}{6}}

\small =-3\pm1.3722\times2.7384\times0.5774

\small =-3\pm2.1695=(-5.1695,-0.8305)

80% confidence interval for µ1 - µ2 = (-5.1695,-0.8305)

left endpoint : -5.1695

right endpoint : -0.8305